# Neural network learns low-dimensional polynomials with SGD near the information-theoretic limit

### v1

@@ A.2 Proof for Non-even Functions $(ii)$
$\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\bigg{(}\frac{g(t)}{2\tau(g)}\bigg{)}^{i}t\bigg{]}.$ (A.13)
@@
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[2\tau(g)\leq|g(t)|<3\tau(g)]\bigg{(}\frac{g(t)}{3\tau(f)}\bigg{)}^{i}t\bigg{]}$ (A.15)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|g(t)|<2\tau(g)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}.$ (A.16)
@@
$\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[|g(t)|\geq 3\tau(f)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.17)
$=\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[\beta\leq t\leq-\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[t\geq-\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.18)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[t\leq\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.19)
$=\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\mathbbm{1}[\beta\leq t\leq-\alpha]t\big{]}+\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\bigg{[}\mathbbm{1}[t\geq-\alpha]\bigg{(}\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}-\bigg{(}\frac{g(-t)}{3\tau(g)}\bigg{)}^{i}\bigg{)}t\bigg{]}$ (A.20)
$>\beta\mathbb{P}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\beta\leq t\leq-\alpha\big{]}.$ (A.21)
@@
$\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\big{]}\geq-\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]|t|\big{]}\bigg{(}\frac{2}{3}\bigg{)}^{i}.$ (A.22)
@@
(A.16)$>\beta\mathbb{P}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\beta\leq t\leq-\alpha\big{]}-\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]|t|\big{]}\bigg{(}\frac{2}{3}\bigg{)}^{i}.$ (A.23)
@@
(ii-3): Lower bound of test function via compactness. Let $\mathcal{F}_{q}$ be a set of unit $L^{2}$-norm polynomials with degree up to $q$ and $\mathbb{E}_{t\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{\mathrm{odd}}(t)^{2}]\geq c$. Since $\mathscr{H}(f)$ is always positive for $\mathcal{F}_{q}$, $\mathscr{H}(f)$ is continuous with respect to $f$, and $\mathcal{F}_{q}$ is a compact set, $\inf_{f\in\mathcal{F}_{q}}\mathscr{H}(f)$ has the minimum value $\mathscr{H}_{0}$ that is positive. Note that $\mathscr{H}(f)$ might depends on $c$.
@@ Appendix B SGD with Reused Batch
In this section we establish the statistical and computational complexity of Algorithm 1. Recall that the algorithm first trains the first-layer parameters with $T_{1}$ steps of SGD update, where we reuse the same sample for two consecutive steps. The analysis of first-layer training is divided into two phases: $(i)$ weak recovery ($\boldsymbol{w}^{\top}\boldsymbol{\theta}\geq\varepsilon$), and $(ii)$ strong recovery ($\boldsymbol{w}^{\top}\boldsymbol{\theta}\geq 1-\varepsilon$). We then train the second-layer parameters after strong recovery is achieved.
The section is organized as follows.
  • Section B.1 verifies the conditions on the activation function $\sigma$ to guarantee weak and strong recovery.
  • Section B.2 isolates a (nearly) constant fraction of neurons at initialization with an alignment $\boldsymbol{w}^{\top}\boldsymbol{\theta}$ above a certain threshold. We focus on such neurons in Phase I of first-layer training.
  • Section B.3 lower bounds the expected update of alignment $\boldsymbol{w}^{\top}\boldsymbol{\theta}$ of two gradient steps, and Section B.4 shows that neurons yield weak recovery within $2T_{1,1}=\tilde{O}(d)$ steps.
  • Section B.5 discusses how to convert weak recovery to strong recovery using $2T_{1,2}=\tilde{O}(d\varepsilon^{-2})$ SGD steps.
  • Finally, Section B.6 analyzes the second-layer training and concludes the proof.
@@
In the following proofs, we introduce constants $c_{i}$ and $C_{i}$, which depends on $d$ at most polylogarithmically. Specifically, the asymptotic strength of the constants is ordered as follows.
$1\simeq C_{1}\lesssim c_{1}^{-1}\lesssim C_{2}\lesssim C_{3}\lesssim c_{2}^{-1}$ (B.1)
@@
$c_{\eta}$ in the main text can be taken as $c_{\eta}=c_{1}$, where $c_{1}$ should satisfy $\lim_{d\to\infty}c_{1}=0$, but the convergence can be arbitrarily slow. This requirement comes from the fact that we do not know the exact value of $H(\sigma_{*}^{I};k_{*})$, which might be very small. To ensure that the signal is isolated, taking $\eta\asymp c_{1}d^{-1}$ with arbitrarily slow $c_{1}$ suffices. $C_{2}$ can also be arbitrarily slow, as long as it satisfies $C_{2}=\mathrm{poly}(c_{1}^{-1})$. $C_{3}=\mathrm{polylog}(d)$ will be used to represent polylogarithmic factor that comes from high probability bounds.
B.1 Conditions on the Activation Function
B.1.1 Verifying Assumption 2
In the following, we focus on the activation function of a single neuron and omit the subscript that distinguishes different neurons. Recall that we consider polynomial activation functions written as
$\sigma(z):=\sum_{j=0}^{C_{\sigma}}\beta_{j}\mathsf{He}_{j}(z).$ (B.2)
For weak recovery, we can use any polynomial that has degree $C_{\sigma}\geq C_{q}$ as long as the following condition holds: If $\mathrm{IE}(\sigma_{*})\geq 2$ and there exists some $i\leq C_{\sigma}$ such that $\mathrm{IE}(\sigma_{*}^{i})=1$, $\sigma$ should satisfy
$H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};0\bigg{)}\neq 0.$ (B.3)
If $\mathrm{IE}(\sigma_{*})\geq 3$ and there does not exist any $i\leq C_{\sigma}$ such that $\mathrm{IE}(\sigma_{*}^{i})=1$ (in this case there exists some $i\leq C_{q}$ such that $\mathrm{IE}(\sigma_{*}^{i})=2$), $\sigma$ should satisfy
$H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};1\bigg{)}\neq 0.$ (B.4)
Below we prove Lemma 2 which shows that the above conditions are met with probability $1$ for randomly drawn the Hermite coefficients.
Proof of Lemma 2. We note that $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)=\mathbb{E}[\sigma^{(i)}(\sigma^{(1)})^{i-1}\mathsf{He}_{k}]$ is a polynomial of $\{\beta_{j}\}_{j=0}^{C_{\sigma}}$. This polynomial is not identically equal to zero. To confirm this, consider $\sigma=x^{C_{\sigma}}+x^{C_{\sigma}-1}$. Because $\sigma^{(i)}(\sigma^{(1)})^{i-1}$ is expanded as a sum of $x^{l}(i(C_{\sigma}-3)\leq l\leq i(C_{\sigma}-2)+1$ with positive coefficients and each $x^{l}$ is a sum of $\mathsf{He}_{l},\mathsf{He}_{l-2}\cdots$ with positive coefficients, $\sigma^{(i)}(\sigma^{(1)})^{i-1}$ has all positive Hermite coefficients for degree $10,1,\cdots,i(C_{\sigma}-2)+$ 1. If $k\leq i(C_{\sigma}-2)+1$, this choice of $\sigma$ yields $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)>0$, which confirms that $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)$ as a polynomial of $\{\beta_{j}\}_{j=0}^{C_{\sigma}}$ is not identically equal to zero.
Now, the assertion follows from the Schwartz–Zippel Lemma [Sch80], or the fact that zeros of a non-zero polynomial form a measure-zero set. ∎
B.1.2 Verifying Assumption 3
On the other hand, for the strong recovery we require an additional condition on the activation function due to link misspecification, which is also introduced in [BAGJ21, MHWSE23]:
$\sum_{j=p}^{p}j!\alpha_{j}\beta_{j}s^{j}>0\quad\text{for all $s>0$}.$ (B.5)
In order to meet Assumption 2 and (3) simultaneously, we follow [OSSW24] and randomize the activation function. Specifically, the activation function should satisfy

### v2

@@ A.2 Proof for Non-even Functions $(ii)$
$\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\bigg{(}\frac{g(t)}{2\tau(g)}\bigg{)}^{i}t\bigg{]}.$ (A.13)
@@
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[2\tau(g)\leq|g(t)|<3\tau(g)]\bigg{(}\frac{g(t)}{3\tau(f)}\bigg{)}^{i}t\bigg{]}$ (A.15)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[|g(t)|<2\tau(g)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}.$ (A.16)
@@
$\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[|g(t)|\geq 3\tau(f)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.17)
$=\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[\beta\leq t\leq-\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[t\geq-\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.18)
$\quad+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[t\leq\alpha]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\bigg{]}$ (A.19)
$=\mathbb{E}_{t\sim\mathcal{N}(0,1)}\big{[}\mathbbm{1}[\beta\leq t\leq-\alpha]t\big{]}+\mathbb{E}_{t\sim\mathcal{N}(0,1)}\bigg{[}\mathbbm{1}[t\geq-\alpha]\bigg{(}\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}-\bigg{(}\frac{g(-t)}{3\tau(g)}\bigg{)}^{i}\bigg{)}t\bigg{]}$ (A.20)
$>\beta\mathbb{P}_{t\sim\mathcal{N}(0,1)}\big{[}\beta\leq t\leq-\alpha\big{]}.$ (A.21)
@@
$\mathbb{E}_{t\sim\mathcal{N}(0,1)}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]\bigg{(}\frac{g(t)}{3\tau(g)}\bigg{)}^{i}t\big{]}\geq-\mathbb{E}_{t\sim\mathcal{N}(0,1)}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]|t|\big{]}\bigg{(}\frac{2}{3}\bigg{)}^{i}.$ (A.22)
@@
(A.16)$>\beta\mathbb{P}_{t\sim\mathcal{N}(0,1)}\big{[}\beta\leq t\leq-\alpha\big{]}-\mathbb{E}_{t\sim\mathcal{N}(0,1)}\big{[}\mathbbm{1}[|g(t)|<2\tau(g)]|t|\big{]}\bigg{(}\frac{2}{3}\bigg{)}^{i}.$ (A.23)
@@
(ii-3): Lower bound of test function via compactness. Let $\mathcal{F}_{q}$ be a set of unit $L^{2}$-norm polynomials with degree up to $q$ and $\mathbb{E}_{t\sim\mathcal{N}(0,1)}[f^{\mathrm{odd}}(t)^{2}]\geq c$. Since $\mathscr{H}(f)$ is always positive for $\mathcal{F}_{q}$, $\mathscr{H}(f)$ is continuous with respect to $f$, and $\mathcal{F}_{q}$ is a compact set, $\inf_{f\in\mathcal{F}_{q}}\mathscr{H}(f)$ has the minimum value $\mathscr{H}_{0}$ that is positive. Note that $\mathscr{H}(f)$ might depends on $c$.
@@
A.3 Proof for Non-Polynomial Functions
For non-polynomial link functions, we note that similar to [DPVLB24], the existence of polynomial basis is needed to exclude extreme cases, and we cannot upper bound the required degree $I$ because general link functions are not included in a compact space.
Proof of Lemma 8. The derivation is analogous to [DPVLB24, Lemma F.14]. Let $z\sim\mathcal{N}(0,1)$ and $y=\sigma_{*}(z)$. We define $\zeta_{p_{*}}(y)=\mathbb{E}[\frac{1}{\sqrt{p_{*}!}}\mathsf{He}_{p_{*}}(z)|y]$ and its basis expansion $\zeta_{p_{*}}(y)=\sum_{k=0}^{\infty}v_{k}\phi_{k}$. Let $K$ be a smallest integer such that $v_{k}\neq 0$. Then, there exists an integer with $I\leq K$ such that $\mathrm{IE}(y^{I})=p_{*}$. Indeed,
$\mathbb{E}[\phi_{K}(y)\mathsf{He}_{p_{*}}(z)]$ $=\mathbb{E}_{y}[\Phi_{K}(y)\mathbb{E}_{z|y}[\mathsf{He}_{p_{*}}(z)|y]]$ (A.26)
$=\mathbb{E}_{y}\bigg{[}\Phi_{K}(y)\sum_{k=0}^{K}v_{k}\phi_{k}(y)\bigg{]}=v_{K}\neq 0,$ (A.27)
which means that at least one of $y,y^{2},\cdots,y^{K}$ yields a non-zero $p_{*}$-th Hermite coefficient. ∎
@@ Appendix B SGD with Reused Batch
In this section we show that Algorithm 1 learns single-index models in $\tilde{O}(d^{1\lor(p_{*}-1)})$ samples with high probability. The algorithm trains the first layer for $T_{1}$ SGD steps, where we sample a new data point in every two steps. The first layer training is further divided into two phases: weak recovery ($\boldsymbol{w}^{\top}\boldsymbol{\theta}\gtrsim 1$) and strong recovery ($\|\boldsymbol{w}-\boldsymbol{\theta}\|\lesssim\varepsilon$). Then, we learn the second layer parameters.
Specifically, Section B.2 shows that at initialization, a (nearly) constant fraction of neurons has alignment $\boldsymbol{w}^{\top}\boldsymbol{\theta}$ beyond a certain threshold. We focus on such neurons in the first phase of training. Section B.3 lower bounds the expected update of alignment $\boldsymbol{w}^{\top}\boldsymbol{\theta}$ of two gradient steps, and Section B.4 establishes that the neurons achieve weak recovery within $2T_{1,1}=\tilde{O}(d^{1\lor(p_{*}-1)})$ steps. Section B.5 discusses how to convert weak recovery to strong recovery using $2T_{1,2}=\tilde{O}(d\varepsilon^{-2})$ more steps. We let $T_{1}=2T_{1,1}+2T_{1,2}$. Finally, Section B.6 analyzes second layer training and concludes the proof.
@@
In the following proofs, we use several constants, which depends on $d$ at most at most polylogarithmically. Specifically, asymptotic strength of the constants is ordered as follows.
$1\simeq c_{\sigma}\simeq C_{1}\lesssim\mathopen{}\mathclose{{}\left\{\begin{matrix}c_{\eta}^{-1}\simeq C_{2}\lesssim\mathrm{poly}(c_{\eta}^{-1})\lesssim\mathopen{}\mathclose{{}\left\{\begin{matrix}c_{1}^{-1}\simeq C_{3}\\ c_{2}^{-1}\\ \end{matrix}}\right\}\\ \delta^{-1}\\ \end{matrix}}\right\}$$\lesssim\mathopen{}\mathclose{{}\left\{\begin{matrix}\delta^{-1}\mathrm{poly}(c_{\eta}^{-1})\lesssim c_{\xi}^{-1}\\ \mathrm{poly}(c_{1}^{-1})\lesssim\bar{c}_{\eta}^{-1}\end{matrix}}\right\}$ (B.1)
$\lesssim\mathrm{polylog}(d)=C_{4}.$ (B.2)
@@
Here, $c_{\eta}$ and $\delta$ should satisfy $\lim_{d\to\infty}c_{\eta}=\lim_{d\to\infty}\delta=0$, but the convergence can be arbitrarily slow, (e.g., as slow as $1/\log\log\log\cdots\log d$). This requirement comes from the fact that we do not know the exact value of $H(\sigma_{*}^{I};p_{*})$. To ensure that one signal term (from the Taylor series) is isolated, taking $\eta\asymp d^{-1}$ with a sufficiently small constant is insufficient but $\eta\asymp c_{\eta}d^{-1}$ with arbitrarily slow $c_{\eta}$ suffices. Also, to guarantee that the failure probability is $o_{d}(1)$, we require $\delta$ to be $o_{d}(1)$. $c_{\xi}$ can also decay arbitrarily slowly, as long as it satisfies $c_{\xi}\lesssim\delta\mathrm{poly}(c_{\eta}^{-1})$. $C_{4}=\mathrm{polylog}(d)$ will be used to represent any polylogarithmic factor that comes from high probability bounds.
For the first-layer training, we can reduce the argument into training of one neuron using the correlation loss as follows. At each step, the gradient update (Line 8 of Algorithm 1) is written as
$=\boldsymbol{w}^{t}_{j}-\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\bigg{)}^{2}+2\eta^{t}_{j}\tilde{\nabla}_{\boldsymbol{w}}\bigg{(}y\frac{1}{N}\sum_{j=1}^{N}a_{j}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\bigg{)}$ (B.4)
$=\boldsymbol{w}^{t}_{j}-\frac{2\eta^{t}c_{a}^{2}}{N}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\bigg{)}\big{(}\tilde{\nabla}_{\boldsymbol{w}}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\big{)}+\frac{2\eta^{t}c_{a}}{N}y\big{(}\tilde{\nabla}_{\boldsymbol{w}}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\big{)}.$ (B.5)
While the second term scales with $\eta^{t}c_{a}^{2}N^{-1}$, the third term scales with $\eta^{t}c_{a}N^{-1}$. Thus, by setting $c_{a}$ sufficiently small, we can ignore the interaction between neurons. We will show that the strength of the signal in the direction of $\boldsymbol{\theta}$ is at least $(\kappa_{j}^{t})^{p_{*}-1}\gtrsim d^{-\frac{p_{*}-1}{2}}$ (up to a polylogarithmic factor, and $p_{*}=\mathrm{GE}(\sigma_{*})$). On the other hand, we can easily see that $\boldsymbol{\theta}^{\top}\big{(}\frac{1}{N}\sum_{j=1}^{N}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\big{)}\big{(}\tilde{\nabla}_{\boldsymbol{w}}\sigma_{j}({\boldsymbol{w}_{j}^{t}}^{\top}\boldsymbol{x})\big{)}$ is bounded by $\tilde{O}(1)$ with high probability. Therefore, by simply letting $c_{a}=\tilde{\Theta}(d^{-\frac{p_{*}-1}{2}})$, we can ignore the effect of the second term in (B.5). Moreover, for simplicity, we will reparameterize $\frac{2\eta^{t}c_{a}}{N}$ as $\eta^{t}$ below. Consequently, we may analyze the following update
$\boldsymbol{w}^{t+1}_{j}\leftarrow\boldsymbol{w}^{t}_{j}+\eta^{t}\tilde{\nabla}_{\boldsymbol{w}}\big{(}y\sigma_{j}({\boldsymbol{w}^{t}_{j}}^{\top}\boldsymbol{x})\big{)},$ (B.6)
instead of Line 8 of Algorithm 1. Since there is no interaction between neurons now, we omit the subscript $j$ when the context is clear.
B.1 Assumptions on Link Function
The analysis consists of three different phases: weak recovery and strong recovery of the first-layer weights, and approximation of the link function (ridge regression of the second-layer). Each phase requires different assumptions on the activation functions, depending on the link function. Before starting the analysis, we decompose Assumptions 2 and 3 and clarify which conditions are needed in each phase. We prove that instead of using a specific activation function tailored to different link functions, a randomized activation function satisfies all required assumptions with probability $\Omega(1)$.
In the following, we write the student activation function as
$\sigma_{j}(s):=\sum_{i=0}^{\infty}\beta_{j,i}\mathsf{He}_{i}(s)$ (B.7)
with coefficients $\{\beta_{j,i}\}_{i=0}^{C_{\sigma}}$ (sometimes the subscript $j$, which is the index of the neurons, is omitted).
B.1.1 For polynomial link functions
In the following, we summarize the precise conditions to be satisfied by the activation functions (these conditions are weaker than Assumptions 2 and 3). For polynomial link functions, we focus on polynomial activation functions (with bounded degree) for simplicity, but non-polynomial activation functions would not change the proof significantly.
Let $p$ and $q$ be the minimum and maximum degree of non-zero Hermite coefficients of $\sigma_{*}$. Note that $\mathrm{GE}(\sigma_{*})=1$ or $2$ holds (see Proposition 6). Let $I\leq C_{q}$ (according to Proposition 6) be the smallest integer such that $\mathrm{IE}(\sigma_{*}^{I})=\mathrm{GE}(\sigma_{*})=p_{*}$ and $C_{\sigma}$ be the degree of the activation function.